# VideoUFO: A Million-Scale User-Focused Dataset for Text-to-Video Generation

### v1

@@ 4.1 Fundamental Attributes
  • VideoUFO is released under a more flexible license and introduces new data. Specifically, we search for videos on YouTube by ourselves and only select those with a Creative Commons license. In contrast, most recent datasets (including Panda-70M [5], LVD-2M [40], Koala-36M [33], VidGen-1M [29], and OpenVid-1M [20]) directly source data from HD-VILA-100M [41]. As a result: (1) they do not contribute new data but rather introduce a new data processing pipeline; (2) they must adhere to the same license as HD-VILA-100M, i.e., the Research Use of Data Agreement (R-UDA), which restricts commercial use. In addition, one of the most widely used datasets, WebVid-10M [3], has been retracted due to potential copyright infringement.
@@ 4.2 Topics Coverage
In this section, we analyze the differences in topics coverage between the proposed VideoUFO and existing datasets.
@@
  • Match the extracted topics with users’ focused ones. We observe that the same or similar topic may be described using different words. For example, both cathedral and church refer to a similar topic. Therefore, when comparing user-focused topics with those from each existing dataset, we choose semantic matching rather than a word-to-word approach. Specifically, (1) we first use $\mathtt{SentenceTransformers}$ [26] to embed the two lists of topics; and (2) then for each user-focused topic, if there exists a topic in the existing dataset with a cosine similarity greater than $0.6$, we consider that user-focused topic to be covered by the existing dataset. Note that the threshold of $0.6$ is an empirical value, as we observe that most similar topics exhibit a cosine similarity greater than $0.6$.
@@
  • None of the existing datasets cover all user-focused topics in VideoUFO. Specifically, the most comprehensive dataset, InternVid [38], covers $821$ topics, accounting for $63.6\%$ of all user-focused topics. Furthermore, if we combine all existing datasets in Table 1 into the “largest” one, the coverage of user-focused topics will increase to $79.5\%$. However, this will lead to a higher computational burden when training text-to-video models.
  • Interestingly, a dataset may cover more user-focused topics than the actual number of topics it contains. For instance, OpenVid-1M [20] contains $671$ topics but covers $703$ user-focused ones. This is reasonable because two topics in our dataset, VideoUFO, may be similar and thus mapped to a single topic in OpenVid-1M [20]. Nevertheless, this does not compromise the quality of our VideoUFO since it only leads to some relative duplicated topics, while the videos in VideoUFO still accurately reflect real users’ focus and interests.
@@ 5.1 Benchmark
Construction. As shown in Fig. 7, the calculation process of our BenchUFO includes: (1) Selecting prompts. First, we select 791 concrete nouns from $1,291$ user-focused topics. (Please refer to the next section for the rationale behind selecting concrete nouns.) Then, for each chosen topic, we randomly select $10$ text prompts from VidProM [35]. All these prompts constitute the prompt set for our benchmark. (2) Generating videos. For each prompt, we use a text-to-video model to generate a corresponding video. (3) Describing videos. For each video, we use a multimodal large language model (here, we choose $\mathtt{QWen2}$-$\mathtt{VL}$-$\mathtt{7B}$ [32]) to understand and describe it. The instruction prompt is provided in the Supplementary (Section 10). (4) Calculating similarity. We use a sentence embedding model (in this case, $\mathtt{SentenceTransformers}$ [26]) to encode both the input prompt and the output description, and then compute the cosine similarity between them. We calculate the average similarity across $10$ prompts for each topic. A higher similarity score indicates better performance on that specific topic. Our benchmark computes and considers the $10$/$50$ worst-performing and $10$/$50$ best-performing topics.
Table 2: The performance of both publicly available text-to-video models and our trained models on the proposed BenchUFO. The publicly available models are trained on various datasets, including both public and private ones. MVDiT [20] is trained on VidGen [29], OpenVid [20], and VideoUFO, respectively. “Low/Top $N$” denotes the average score of the worst/best-performing $N$ topics.
@@
Mira [10] $0.236$ $0.282$ $0.508$ $0.550$
Show-1 [44] $0.266$ $0.303$ $0.524$ $0.564$
LTX-Video [15] $0.268$ $0.310$ $0.532$ $0.574$
Open-Sora-Plan [16] $0.314$ $0.361$ $0.559$ $0.598$
TF-T2V [36] $0.316$ $0.359$ $0.560$ $0.595$
Mochi-1 [30] $0.323$ $0.367$ $0.580$ $0.616$
HiGen [24] $0.352$ $0.394$ $0.589$ $0.625$
Open-Sora [45] $0.363$ $0.409$ $0.601$ $0.639$
Pika [22] $0.365$ $0.404$ $0.583$ $0.619$
RepVideo [28] $0.368$ $0.402$ $0.589$ $0.619$
T2V-Zero [11] $0.375$ $0.410$ $0.586$ $0.616$
CogVideoX [42] $0.383$ $0.419$ $0.601$ $0.629$
Latte-1 [18] $0.384$ $0.421$ $0.592$ $0.636$
HunyuanVideo [12] $0.388$ $0.427$ $0.612$ $0.645$
LaVie [37] $0.399$ $0.426$ $0.595$ $0.632$
Pyramidal [8] $0.400$ $0.433$ $0.606$ $0.647$
MVDiT-VidGen [29] $0.382$ $0.426$ $0.594$ $0.626$
MVDiT-OpenVid [20] $0.401$ $0.437$ $0.609$ $0.645$
@@
  • Why use video understanding model reasonable? This is because the video understanding model has advanced significantly with the advent of large language models, and it is reasonable to assume that these models have been exposed to images/videos containing user-focused topics. Here, we choose $\mathtt{QWen2}$-$\mathtt{VL}$-$\mathtt{7B}$ [32] as the video understanding model. Trained on $800$ billion tokens of visual-related data and $600$ billion tokens of text data, it is expected to accurately comprehend these user-focused topics and faithfully summarize a video.
  • Why not use established benchmarks? This is because the existing benchmarks fail to reflect real-world scenarios.. Specifically, we note that there exists a well-established benchmark, VBench [7], whose “object class” and “multiple objects” dimensions are similar to our BenchUFO. However, the prompts in VBench have two main drawbacks: (1) They cover only a limited number of topics/objects – specifically, just $79$. (2) They focus exclusively on common topics/objects (e.g., “bicycle”, “car”, and “airplane”), while many topics that users care about, such as “forest”, “sunset”, and “beach”, are missing.
Figure 8: Visual comparisons between our approach (MVDiT-VideoUFO) and other text-to-video models. The model trained on VideoUFO outperforms the alternatives in generating user-focused topics. In the prompts, concepts in green indicate those successfully generated solely by our method, whereas those in yellow denote concepts successfully generated by both our model and the competing models. Here we represent each video with a single frame, and all generated videos are provided in the supplementary materials.
@@ 5.2 Observations
  • The proposed VideoUFO helps reduce the gap between the worst-performing and best-performing topics. To demonstrate the effectiveness of the proposed VideoUFO, we train an MVDiT [20] solely on VideoUFO. We find that the trained model achieves the highest low-$10$ scores (a $+4.2\%$ improvement compared to the current state-of-the-art) while maintaining performance on the top-$10$ topics. Visually, the trained model successfully generates videos on topics that other models previously could not.
  • The proposed VideoUFO outperforms other similar-scale datasets, such as OpenVid [20] and VidGen [29]. To demonstrate that the improvement originates from our VideoUFO rather than from the MVDiT architecture [20], we replace our VideoUFO with OpenVid [20] and VidGen [29]. We find that when using the popular OpenVid [20], the low-$10$ scores are similar to previous state-of-the-art models ($0.401$ vs. $0.400$). Furthermore, when training on VidGen [29], the performance is even lower.
5.3 Future works
Due to limited computational resources, when validating the effectiveness of our dataset, we currently follow the video generation method from the academic community introduced by [20], which already requires $32$ A100 GPUs. However, the quality of videos generated by this baseline remains limited. In the future, industry researchers could incorporate our dataset into their large-scale foundational model training pipelines to further explore its potential.
@@ 6 Ablation Studies on Curation of VideoUFO
This section performs ablation studies to investigate key components in curating VideoUFO, i.e., examining (1) the impact of the number of clusters in user-focused topic analysis and (2) the methods for clip verification. For more details, please refer to Supplementary (Section 11).
@@ Acknowledgment
Supplementary Material
8 The Prompt for Concluding Topics
@@
9 The Distributions of Video Quality Score
10 The Prompt for Describing Videos
@@
11 Ablation Studies on Curation of VideoUFO

### v2

@@ 4.1 Fundamental Attributes
  • VideoUFO is released under a more flexible license and introduces new data. Specifically, we search for videos on YouTube by ourselves and only select those with a Creative Commons license. In contrast, most recent datasets (including Panda-70M chen2024panda , LVD-2M xiong2024lvd , Koala-36M wang2024koala , VidGen-1M tan2024vidgen , and OpenVid-1M nan2024openvid ) directly source data from HD-VILA-100M xue2022advancing . As a result: (1) they do not contribute new data but rather introduce a new data processing pipeline; (2) they must adhere to the same license as HD-VILA-100M, i.e., the Research Use of Data Agreement (R-UDA), which restricts commercial use. In addition, one of the most widely used datasets, WebVid-10M bain2021frozen , has been retracted due to potential copyright infringement.
@@ 4.2 Topics Coverage
This section analyzes the differences in topics coverage between the VideoUFO and existing datasets.
@@
  • Match the extracted topics with users’ focused ones. We observe that the same or similar topic may be described using different words. For example, both cathedral and church refer to a similar topic. Therefore, when comparing user-focused topics with those from each existing dataset, we choose semantic matching rather than a word-to-word approach. Specifically, (1) we first use $\mathtt{SentenceTransformers}$ reimers2019sentencebert to embed the two lists of topics; and (2) then for each user-focused topic, if there exists a topic in the existing dataset with a cosine similarity greater than $0.6$, we consider that user-focused topic to be covered by the existing dataset. Note that the threshold of $0.6$ is an empirical value, as we observe that most similar topics exhibit a cosine similarity greater than $0.6$.
@@
  • None of the existing datasets cover all user-focused topics in VideoUFO. Specifically, the most comprehensive dataset, InternVid wanginternvid , covers $821$ topics, accounting for $63.6\%$ of all user-focused topics. Furthermore, if we combine all existing datasets in Table 1 into the “largest” one, the coverage of user-focused topics will increase to $79.5\%$. However, this will lead to a higher computational burden when training text-to-video models.
  • Interestingly, a dataset may cover more user-focused topics than the actual number of topics it contains. For instance, OpenVid-1M nan2024openvid contains $671$ topics but covers $703$ user-focused ones. This is reasonable because two topics in our dataset, VideoUFO, may be similar and thus mapped to a single topic in OpenVid-1M nan2024openvid . Nevertheless, this does not compromise the quality of our VideoUFO since it only leads to some relative duplicated topics, while the videos in VideoUFO still accurately reflect real users’ focus and interests.
@@ 5.1 Benchmark
Construction. As shown in Fig. 6, the calculation process of our BenchUFO includes: (1) Selecting prompts. First, we select 791 concrete nouns from $1,291$ user-focused topics. (Please refer to the next section for the rationale behind selecting concrete nouns.) Then, for each chosen topic, we randomly select $10$ text prompts from VidProM wang2024vidprom . All these prompts constitute the prompt set for our benchmark. (2) Generating videos. For each prompt, we use a text-to-video model to generate a corresponding video. (3) Describing videos. For each video, we use a multimodal large language model (here, we choose $\mathtt{QWen2}$-$\mathtt{VL}$-$\mathtt{7B}$ Qwen2VL ) to understand and describe it. The instruction prompt is provided in the Appendix (Section B). (4) Calculating similarity. We use a sentence embedding model (in this case, $\mathtt{SentenceTransformers}$ reimers2019sentencebert ) to encode both the input prompt and the output description, and then compute the cosine similarity between them. We calculate the average similarity across $10$ prompts for each topic. A higher similarity score indicates better performance on that specific topic. Our benchmark computes and considers the $10$/$50$ worst-performing and $10$/$50$ best-performing topics.
  • Why use video understanding model reasonable? This is because the video understanding model has advanced significantly with the advent of large language models, and it is reasonable to assume that these models have been exposed to images/videos containing user-focused topics. Here, we choose $\mathtt{QWen2}$-$\mathtt{VL}$-$\mathtt{7B}$ Qwen2VL as the video understanding model. Trained on $800$ billion tokens of visual-related data and $600$ billion tokens of text data, it is expected to accurately comprehend these user-focused topics and faithfully summarize a video.
  • Why not use established benchmarks? This is because the existing benchmarks fail to reflect real-world scenarios.. Specifically, we note that there exists a well-established benchmark, VBench huang2024vbench , whose “object class” and “multiple objects” dimensions are similar to our BenchUFO. However, the prompts in VBench have two main drawbacks: (1) They cover only a limited number of topics/objects – specifically, just $79$. (2) They focus exclusively on common topics/objects (e.g., “bicycle”, “car”, and “airplane”), while many topics that users care about, such as “forest”, “sunset”, and “beach”, are missing.
Figure 7: The performance of both publicly available text-to-video models and our trained models on the proposed BenchUFO. The publicly available models are trained on various datasets, including both public and private ones. MVDiT nan2024openvid is trained on VidGen tan2024vidgen , OpenVid nan2024openvid , and VideoUFO, respectively. “Low/Top $N$” denotes the average score of the worst/best-performing $N$ topics.
@@
Mira jumiradata $0.236$ $0.282$ $0.508$ $0.550$
Show-1 zhang2024show $0.266$ $0.303$ $0.524$ $0.564$
LTX-Video ltxvideo $0.268$ $0.310$ $0.532$ $0.574$
Open-Sora-Plan lin2024open $0.314$ $0.361$ $0.559$ $0.598$
TF-T2V wang2024recipe $0.316$ $0.359$ $0.560$ $0.595$
Mochi-1 genmo2024mochi $0.323$ $0.367$ $0.580$ $0.616$
HiGen qing2024hierarchical $0.352$ $0.394$ $0.589$ $0.625$
Open-Sora opensora $0.363$ $0.409$ $0.601$ $0.639$
Pika PikaLabs $0.365$ $0.404$ $0.583$ $0.619$
RepVideo si2025repvideo $0.368$ $0.402$ $0.589$ $0.619$
T2V-Zero khachatryan2023text2video $0.375$ $0.410$ $0.586$ $0.616$
CogVideoX yang2024cogvideox $0.383$ $0.419$ $0.601$ $0.629$
Latte-1 ma2024latte $0.384$ $0.421$ $0.592$ $0.636$
HunyuanVideo kong2024hunyuanvideo $0.388$ $0.427$ $0.612$ $0.645$
LaVie wang2024lavie $0.399$ $0.426$ $0.595$ $0.632$
Pyramidal jin2024pyramidal $0.400$ $0.433$ $0.606$ $0.647$
MVDiT-VidGen tan2024vidgen $0.382$ $0.426$ $0.594$ $0.626$
MVDiT-OpenVid nan2024openvid $0.401$ $0.437$ $0.609$ $0.645$
@@ 5.2 Observations
  • The proposed VideoUFO helps reduce the gap between the worst-performing and best-performing topics. To demonstrate the effectiveness of the proposed VideoUFO, we train an MVDiT nan2024openvid solely on VideoUFO. We find that the trained model achieves the highest low-$10$ scores (a $+4.2\%$ improvement compared to the current state-of-the-art) while maintaining performance on the top-$10$ topics. Visually, the trained model successfully generates videos on topics that other models previously could not.
  • The proposed VideoUFO outperforms other similar-scale datasets, such as OpenVid nan2024openvid and VidGen tan2024vidgen . To demonstrate that the improvement originates from our VideoUFO rather than from the MVDiT architecture nan2024openvid , we replace our VideoUFO with OpenVid nan2024openvid and VidGen tan2024vidgen . We find that when using the popular OpenVid nan2024openvid , the low-$10$ scores are similar to previous state-of-the-art models ($0.401$ vs. $0.400$). Furthermore, when training on VidGen tan2024vidgen , the performance is even lower.
Limitation. Due to limited computational resources, when validating the effectiveness of our dataset, we currently follow the video generation method from the academic community introduced by nan2024openvid , which already requires $32$ A100 GPUs. However, the quality of videos generated by this baseline remains limited. In the future, industry researchers could incorporate our dataset into their large-scale foundational model training pipelines to further explore its potential.
Figure 8: Visual comparisons between our approach (MVDiT-VideoUFO) and other text-to-video models. The model trained on VideoUFO outperforms the alternatives in generating user-focused topics. In the prompts, concepts in green indicate those successfully generated solely by our method, whereas those in yellow denote concepts successfully generated by both our model and the competing models. All generated videos are provided in the supplementary materials.
@@ 6 Ablation Studies on Curation of VideoUFO
This section performs ablation studies to investigate key components in curating VideoUFO, i.e., examining (1) the impact of the number of clusters in user-focused topic analysis and (2) the methods for clip verification. For more details, please refer to Appendix (Section C).
@@ Acknowledgments
Appendix A The Prompt for Concluding Topics
@@
Appendix B The Prompt for Describing Videos
@@
Appendix C Ablation Studies on Curation of VideoUFO